ecītes it to SO atacante The value, r, ofthe equiment is a finction ofthe number ofyears, L Asmall business has spent S32.500 on new compde pennt-nd, fix lax purposes, At r 6 years from the date of parchase, its value is $2500 The domain of Pt) is hinter İsts The range of PtO is
Answer #1

Let the function be V(t) = 32500-mt.

Given, when t = 6, then V(t) = 2500.

Putting this value in the above function we get,

2500 = 32500-6m

i.e., 6m = 32500-2500

i.e., m = 30000/6

i.e., m = 5000

Then, the function is : V(t) = 32500-5000t.

Therefore, the domain of V(t) is : 0<t<6 .

And, the range of V(t) is : 2500\leq t\leq32500 .